OpenStudy (anonymous):

What is the range of f(x) = 1 - √x

OpenStudy (anonymous):

Let f(x) = y So: y = 1 - root(x) So find x in terms of y : root(x) = 1 - y Square both the sides: x = (1 - y )^2 Clearly, y can take all the Real Values..

OpenStudy (anonymous):

By plotting the function also, u can find the range:)

OpenStudy (anonymous):

\[\sqrt{x}\ge0\]\[-\sqrt{x}\le0\]\[1-\sqrt{x}\le1\]